# Scalable In-Context Q-Learning

### v1

@@ 4.1 Prompt Construction via World Modeling
The environment dynamics, i.e., the reward and state transition functions, can share some common structure across the task distribution. For each task $M^{i}$, we approximate its dynamics by a generalized world model $P$ that is shared across tasks, defined as
@@
As the true task identity is unknown, we infer task representation $z_{t}^{i}$ from the agent’s $k$-step experience within task $M^{i}$ as $\eta_{t}^{i}\!=\!(s_{t-k},a_{t-k},r_{t-k},...,s_{t-1},a_{t-1},r_{t-1},s_{t},a_{t})^{i}$. The intuition is that the true task belief can be inferred from the agent’s history interactions, similar to recent meta-RL studies (zintgraf2021varibad, ; ni2023metadiffuser, ). We use a context encoder $E_{\phi}$ to abstract recent $k$-step experiences into a task representation as $z^{i}_{t}\!=\!E_{\phi}(\eta_{t}^{i})$, which is augmented into the input of a dynamics decoder $D_{\varphi}$ to predict the instant reward and next state as $[\hat{r}_{t},\hat{s}_{t+1}]\!=\!D_{\varphi}(s_{t},a_{t};z_{t}^{i})$. Based on the assumption that tasks with similar contexts will behave similarly (lee2020context, ), our generalized world model can extrapolate meta-level knowledge across tasks by precisely capturing task-relevant information. The context encoder and dynamics decoder are jointly trained by minimizing the reward and state transition prediction error as
$\mathcal{L}(\phi,\varphi)=\mathbb{E}_{\eta_{t}^{i}\sim M_{i}}\!\left[\mathbb{E}_{z^{i}_{t}\sim E_{\phi}(\eta_{t}^{i})}\!\left[\lVert[r_{t},s_{t+1}]-D_{\varphi}(s_{t},a_{t};z_{t}^{i})\rVert^{2}\right]\right],~{}~{}~{}\forall~{}\text{task}~{}i.$ (2)
After proper pretraining, we freeze the generalized world model for prompt construction during policy training. At each training step, we sample a short $h$-step trajectory $[s_{1},a_{1},r_{1},...,s_{h},a_{h},r_{h}]^{i}$ randomly from dataset $\mathcal{D}^{i}$, and use the pretrained context encoder $E_{\phi}$ to transform raw transitions in this trajectory into compact task representations as
$\beta^{i}:=[z_{1},...,z_{h}]^{i}=E_{\phi}\left([s_{1},a_{1},r_{1},...,s_{h},a_{h},r_{h}]^{i}\right),~{}~{}~{}\forall~{}\text{task}~{}i.$ (3)
Then, we construct a lightweight prompt $\beta$ using the $h$-step task representation for fast and precise in-context inference, as opposed to AD algorithms that can require long training histories as context.
@@ 4.2 In-Context Q-Learning
The mainstream ICRL methods adopt a supervised learning routine to train transformer policies. However, they often lack the capacity for policy stitching, one of the fundamental abilities of the full RL paradigm to learn optimal policies from suboptimal data (yamagata2023q, ). This issue becomes more pronounced when the offline dataset consists solely of suboptimal trajectories, especially for pure imitation learning approaches that rely on an oracle for optimal action labeling, such as DPT. Classical RL algorithms learn a value function to backpropagate expected returns using dynamic programming updates, showcasing high stitching capacity that allows for not only leveraging collected data but also achieving substantial improvement beyond it (sutton1998reinforcement, ). Inspired by this, we harness the stitching property to offer a promising avenue for unleashing ICRL’s potential toward explicit reward maximization.
In-Context Value Functions. Following common practice in advanced offline RL algorithms (kostrikov2022offline, ; snell2023offline, ), we learn both a state value function $V$ and an action value function $Q$. To maintain scalability and parameter efficiency, we design a multi-head transformer architecture to predict the policy and value functions using three separate heads. Let $\theta$ denote parameters of the integrated transformer model. The model outputs the policy $\pi_{\theta}(a|s;\beta)$, the state value $V_{\theta}(s;\beta)$, and the action value $Q_{\theta}(s,a;\beta)$, given the task prompt $\beta$ and corresponding query inputs. The in-context Q-function is trained to minimize the Bellman bootstrapping error as
@@
where $L_{2}^{\omega}(u)=|\omega-\mathbbm{1}(u<0)|\cdot u^{2}$ is an asymmetric loss function with a expectile parameter $\omega\in(0,1)$, introduced in implicit Q-learning (kostrikov2022offline, ). This form of expectile regression reduces the influence of $Q\!<\!V$ predictions by a factor of $1\!-\!\omega$ while assigning more importance to $Q\!>\!V$ predictions by a factor of $\omega$. In this way, we predict an upper-expectile of the temporal-difference target that approximates the maximum of $r(s_{t}^{i},a_{t}^{i})+\gamma Q_{\hat{\theta}}(s_{t+1}^{i},a_{t+1}^{i};\beta^{i})$ over actions $a_{t+1}^{i}$ constrained to the dataset actions. The algorithm alternates between fitting the state value function with expectile regression and using it to compute Bellman backups for training the Q-function, performing efficient dynamic programming updates for multiple iterations.
In-Context Policy Extraction. The value function learning procedure learns an in-context approximation of optimal Q-functions across tasks, allowing for stitching suboptimal trajectories to discover globally optimal behaviors through value bootstrapping. Then, we distill the in-context value functions into policy extraction with advantage-weighted regression (peng2019advantage, ), a supervised learning style that uses a simple and convergent maximum likelihood loss function as
@@
where $\lambda\!>\!0$ is a temperature parameter. Using weights derived from in-context value functions, the objective is not merely to clone all behaviors from the dataset, but to learn a policy that maximizes the Q-values subject to a distribution constraint imposed by dataset actions. This formulation allows us to select and stitch the most optimal actions in the dataset, while retaining the scalability and stability of the supervised pretraining paradigm.
Finally, the backbone and three separate heads are jointly updated by the overall loss function as
@@ 5 Experiments
We comprehensively evaluate the in-context learning capacity of SICQL on popular benchmarking domains across different dataset types. In general, we aim to answer the following questions:
  • Can SICQL achieve consistent performance gain on few-shot generalization ability to unseen tasks compared to other strong baselines? (Sec. 5.1)
@@
  • Is SICQL robust to the quality of offline datasets and hyperparameters? (Sec. 5.3 and Sec. 5.4)
Environments. We conduct experiments on three challenging benchmarks commonly used for evaluating meta-RL and ICRL algorithms with a multi-task setting: i) DarkRoom laskin2022context , a 2D discrete environment where the agent must locate an unknown goal within a $10\!\times\!10$ room, receiving a reward of 1 only when it reaches the goal. ii) The multi-task MuJoCo control todorov2012mujoco , a well-established testbed that contains two environments where tasks differ in reward functions (Cheetah-Vel and Ant-Dir) and one environment where tasks differ in transition dynamics (Walker-Param). iii) The Meta-World ML1 benchmark suite yu2020meta , a popular robotic platform where a simulated Sawyer arm is designed to perform a wide range of $50$ robotic manipulation tasks. We consider two environments of Reach-v2 and Push-v2, where the robotics arm needs to reach a randomly assigned goal position in 3D space and push a block to a randomly placed target on a tabletop. For each environment, we sample tasks at random from the task distribution and split them into a training set $M^{\text{train}}$ and a test set $M^{\text{test}}$. More details about evaluation environments are provided in Appendix C.
Pretraining Datasets. For DarkRoom, we use a noisy action selection strategy during dataset collection: with probability $\epsilon$, the optimal policy is applied, and with probability $1\!-\!\epsilon$, a random policy is used. The value of $\epsilon$ is set to make sure that the average return in the pretraining dataset is less than 40% of the optimum, thus constructing a suboptimal dataset. We use Soft Actor-Critic (SAC) (haarnoja2018soft, ) for MuJoCo and Proximal Policy Optimization (PPO) (schulman2017proximal, ) for Meta-World to train a single-task policy independently for collecting the dataset of each task. We construct two qualities of offline datasets: Mixed and Medium. More details on dataset construction are provided in Appendix D.
Baselines. We compare to six competitive ICRL approaches and one offline meta-RL method, including: 1) IC-IQL (tarasov2025yes, ), 2) DIT (dong2024incontext, ), 3) DICP (son2025distilling, ), 4) IDT (huang2024context, ), 5) AD (laskin2022context, ), 6) DPT (lee2023supervised, ), and 7) UNICORN (li2024towards, ). Details about these baselines are provided in Appendix E.
To ensure a fair comparison, we conduct a few-shot evaluation for all methods. During testing, the policy directly interacts with the environment for a few episodes with fixed policy parameters, conditioned on a prompt sampled from history interactions. The reported results represent the mean of 10 evaluation trials, with mean returns accompanied by 95% bootstrapped confidence intervals. Standard errors are also provided for the numerical results.
Figure 2: Few-shot evaluation return curves of SICQL and baselines on Mixed datasets.
Table 1: Few-shot evaluation returns of SICQL and baselines on Mixed datasets, i.e., numerical results of converged performance from Figure 2.
@@ 5.1 Main Results
The baselines typically work under the few-shot setting, since they rely on task prompts or warm-start data to infer task representations. We evaluate SICQL against these baselines in an aligned few-shot setting, where all methods utilize the same number of interaction trajectories for task inference. Figure 2 illustrates the evaluation return curves across various environments using Mixed datasets, and Table 1 summarizes the numerical results of converged performance. Across environments with varying reward functions and transition dynamics, SICQL consistently demonstrates superior data efficiency and higher asymptotic performance. In more complex environments such as Ant-Dir and HalfCheetah-Vel, the advantage of SICQL is more pronounced, highlighting its promising in-context learning abilities on challenging tasks. Moreover, SICQL generally exhibits lower variance during the testing phase, indicating both improved data efficiency and better training stability.
@@
Figure 3: Few-shot evaluation return curves of SICQL and its ablations on Mixed datasets. w/o_c removes world modeling, w/o_q removes Q-learning, and w/o_cq removes both components.
Table 2: Few-shot evaluation returns of SICQL and its ablations on Mixed datasets, i.e., numerical results of converged performance from Figure 3.
Environments w/o_cq w/o_c w/o_q SICQL
Reach 736.72 $\pm$ 39.42 792.09 $\pm$ 6.66 752.41 $\pm$ 24.15 806.97 $\pm$ 6.35
Cheetah-Vel -78.35 $\pm$ 4.50 -56.19 $\pm$ 1.94 -63.66 $\pm$ 8.26 -35.48 $\pm$ 1.33
Ant-Dir 591.31 $\pm$ 45.54 693.87 $\pm$ 30.41 784.07 $\pm$ 24.88 813.34 $\pm$ 14.12
@@ 5.2 Ablation Study
To assess the respective contribution of each component, we compare SICQL with three ablations: (i) w/o_c, it removes the world modeling component and directly uses a trajectory of raw transitions to construct prompts; (ii) w/o_q, it removes the Q-learning component and learn in-context policies in a pure supervised paradigm; and (iii) w/o_cq, which removes both world modeling and Q-learning, reducing the model to the original DPT. In all ablations, the remaining structural components are kept identical to those in the full SICQL.
Figure 3 and Table 2 present the ablation results across representative environments. First, removing any individual component significantly degrades SICQL’s performance, highlighting the critical role of each module and their nontrivial integration. Second, ablating the world model leads to a notable performance drop, especially in complex environments like Ant-Dir, which emphasizes the importance of world modeling for precise in-context inference. Third, removing the Q-learning component reduces performance, underscoring its role in leveraging suboptimal data for efficient policy refinement. Integrating Q-learning allows SICQL to improve policies from noisy or inferior trajectories, enhancing the stitching ability for reward maximization. Finally, removing both the world model and Q-learning results in the greatest performance degradation, reducing the model to a baseline form of behavioral cloning that fails to generalize in complex tasks. This confirms the indispensability of both components for robust performance in ICRL. Overall, the ablation result validates SICQL’s effectiveness in harnessing both world modeling and Q-learning for efficient reward maximization and generalization across tasks.
Figure 4: Few-shot evaluation return curves of SICQL and baselines on Medium datasets.
@@ 5.3 Robustness to the Quality of Offline Datasets
Table 3: Few-shot evaluation returns of SICQL and baselines using Medium datasets, i.e., numerical results of converged performance.
@@
UNICORN 366.53$\pm$ 1.98 -103.23$\pm$ 19.35 331.23$\pm$ 25.75
DPT 714.70$\pm$ 32.51 -161.88$\pm$ 12.61 73.72$\pm$ 34.71
AD 626.24$\pm$ 19.08 -77.30$\pm$ 2.06 183.51$\pm$ 11.82
IDT 604.49$\pm$ 3.54 -103.23$\pm$ 19.35 305.50$\pm$ 30.20
DICP 719.54$\pm$ 10.20 -80.34$\pm$ 5.82 416.58$\pm$ 16.80
DIT 709.07$\pm$ 33.65 -82.61$\pm$ 3.15 169.22$\pm$ 5.18
IC-IQL 726.68$\pm$ 15.61 -74.88$\pm$ 2.78 500.20$\pm$ 31.49
SICQL 743.72$\pm$ 22.15 -58.48$\pm$ 2.32 600.18$\pm$ 21.46
To evaluate SICQL’s robustness to data quality, we conduct experiments on Medium datasets that only contain suboptimal data. Figure 4 and Table 3 show SICQL’s performance against baselines on representative environments. SICQL consistently outperforms all baselines, especially in complex environments such as Cheetah-Vel and Ant-Dir, showcasing its superiority when learning from suboptimal trajectories. Notably, many baselines converge to suboptimal policies using imperfect data, especially for pure imitation learning methods like DPT. This again highlights our advantage of using Q-learning to not only simply leverage collected data, but also to combine parts of suboptimal trajectories for finding globally optimal behaviors. Overall, the above results, together with those in Sec. 5.1, confirm the robustness of our method to dataset quality.
5.4 Hyperparameter Analysis
Table 4: Few-shot results of SICQL’s performance with varying $\lambda$ on Mixed datasets.
$\lambda$ Reach Cheetah-Vel Ant-Dir
1.0 765.99$\pm$ 17.93 -99.16$\pm$ 4.00 486.23$\pm$ 30.98
0.1 796.42$\pm$ 7.20 -46.54$\pm$ 2.37 749.00$\pm$ 19.55
0.01 799.29$\pm$ 2.77 -46.56$\pm$ 3.08 813.34$\pm$ 14.12
0.001 806.97$\pm$ 6.35 -35.48$\pm$ 1.33 767.55$\pm$ 19.89
The temperature parameter $\lambda$ in Eq.(6) plays a crucial role in balancing the trade-off between behavior cloning and the greedy pursuit of high Q-values. We conduct experiments to analyze the influence of $\lambda$ on SICQL’s performance. Table 4 present the few-shot ablation results across representative domains with varying values of $\lambda$. A smaller $\lambda$ will make the distribution of advantage weights $\exp(\frac{1}{\lambda})(Q-V)$ less uniform, leading to more exploitation of Q-learning. As $\lambda$ decreases from $1.0$, the performance of SICQL can be greatly improved, highlighting the essentiality of harnessing Q-learning to steer ICRL architectures toward fundamental reward maximization. In practice, a small value of $\lambda=0.01$ or $\lambda=0.001$ leads to a satisfactory performance. The result further confirms the effectiveness and necessity of our in-context Q-learning.
Table 5: Few-shot results of SICQL’s performance with varying $k$ on Mixed datasets.
$K$ Reach Cheetah-Vel Ant-Dir
$2$ 802.00 $\pm$ 10.80 -37.51 $\pm$ 1.80 717.95 $\pm$ 37.81
$4$ 806.97 $\pm$ 6.35 -35.48 $\pm$ 1.33 813.34 $\pm$ 14.12
$6$ 794.19 $\pm$ 7.93 -40.36 $\pm$ 1.36 648.70 $\pm$ 56.75
In prompt construction, the task representation $z_{t}^{i}$ is inferred from the agent’s $k$-step experience $\eta_{t}^{i}$ as $z_{t}^{i}\sim E_{\phi}(\eta_{t}^{i})$. The step $k$ plays a crucial role in capturing task-relevant information. We conduct experiments to investigate the impact of $k$ on SICQL’s performance, and Table 5 presents the few-shot evaluation returns of SICQL across representative environments with varying values of $k$. The results show that the performance of SICQL is not sensitive to $k$, with a moderate experience step of $k=4$ performing the best in all evaluated domains. A small step may provide insufficient task-relevant information, limiting precise in-context inference. A large step may introduce redundancy and noise in the recent experience, and lead to overfitting during world model pretraining. Also, a large step will increase the computation load in pretraining the world model and decrease the speed in constructing the prompt for in-context task inference. In practice, a moderate value of $k$ can achieve both fast and precise in-context inference, highlighting the superiority of our prompt design via world modeling.

### v2

@@ 3.3 World Model Pretraining for Prompt Construction
The environment dynamics, i.e., the reward and state transition functions $p(s^{\prime},r|s,a)$, can share some common structure across the task distribution. For each task $M^{i}$, we approximate its dynamics by a generalized world model $P$ that is shared across tasks, defined as
@@
As the true task identity is unknown, we infer task representation $z_{t}^{i}$ from the agent’s $k$-step experience within task $M^{i}$ as $\eta_{t}^{i}\!=\!(s_{t-k},a_{t-k},r_{t-k},...,s_{t-1},a_{t-1},r_{t-1},s_{t},a_{t})^{i}$. The intuition is that the true task belief can be inferred from the agent’s history interactions, similar to recent meta-RL studies (Zintgraf et al., 2021; Ni et al., 2023). We use a context encoder $E_{\phi}$ to abstract recent $k$-step experiences into a task representation as $z^{i}_{t}\!=\!E_{\phi}(\eta_{t}^{i})$, which is augmented into the input of a dynamics decoder $D_{\varphi}$ to predict the instant reward and next state as $\hat{r}_{t},\hat{s}_{t+1}]\!=\!D_{\varphi}(s_{t},a_{t};z_{t}^{i})$. Based on the assumption that tasks with similar contexts will behave similarly (Lee et al., 2020), our generalized world model can extrapolate meta-level knowledge across tasks by precisely capturing task-relevant information. The context encoder and dynamics decoder are jointly trained by minimizing the reward and state transition prediction error as
$\mathcal{L}(\phi,\varphi)=\mathbb{E}_{\eta_{t}^{i}\sim M^{i}}\!\left[\mathbb{E}_{z^{i}_{t}\sim E_{\phi}(\eta_{t}^{i})}\!\left[\lVert[r_{t},s_{t+1}]-D_{\varphi}(s_{t},a_{t};z_{t}^{i})\rVert^{2}\right]\right],~~~\forall~\text{task}~i.$ (3)
After proper pretraining, we freeze the generalized world model for prompt construction in Eq. (1).
@@ 3.4 In-Context Q-Learning
In-Context Value Functions. The Q-function is trained to minimize the Bellman error as
@@
where $L_{2}^{\omega}(u)=|\omega-\mathbbm{1}(u<0)|\cdot u^{2}$ is an asymmetric loss function with a expectile parameter $\omega\in(0.5,1)$. This form of expectile regression reduces the influence of Q<VQ\!<\!V predictions by a factor of $1\!-\!\omega$ while assigning more importance to Q>VQ\!>\!V predictions by a factor of $\omega$. In this way, we predict an upper-expectile of the temporal-difference target that approximates the maximum of $r(s_{t}^{i},a_{t}^{i})+\gamma Q_{\hat{\theta}}(s_{t+1}^{i},a_{t+1}^{i};\beta^{i})$ over actions $a_{t+1}^{i}$ constrained to the dataset actions. More details can be found in Implicit Q-Learning (Kostrikov et al., 2022).
In-Context Policy Extraction.
The value function learning procedure allows for stitching suboptimal trajectories to discover globally optimal behaviors. Then, we distill the in-context value functions into policy extraction with advantage-weighted regression (Peng et al., 2019), a supervised learning style that uses a simple and convergent maximum likelihood loss function as
@@
where $\lambda\!>\!0$ is a temperature parameter. Using weights from in-context value functions, the objective is not merely to clone behaviors from the dataset but to learn policies that maximize Q-values under a distribution constraint from dataset actions. This formulation aims to select and stitch optimal actions in the dataset while retaining the scalability and stability of the supervised pretraining paradigm.
3.5 Overall Optimization
To unify policy learning and value updating within a single architecture, the transformer backbone and its three specific heads are jointly optimized. This joint training objective combines supervised training with dynamic programming updates, allowing the model to simultaneously learn in-context policies and value functions while retaining scalability and stability. The overall loss is defined as
@@ 3.5 Overall Optimization
This design ensures that S-ICQL integrates lightweight prompt construction with a unified optimization pipeline, achieving efficient reward maximization and robust generalization across tasks.
@@ 4 Experiments
We comprehensively evaluate the in-context learning capacity of S-ICQL on popular benchmarking domains across different dataset types. In general, we aim to answer the following questions:
  • Can S-ICQL achieve consistent performance gain on few-shot generalization ability to unseen tasks compared to other strong baselines? (Sec. 4.1)
@@
  • Can S-ICQL also achieve well performance for out-of-distribution (OOD) tasks? (Sec. 4.3)
  • Does S-ICQL really have the stitching capacity to find globally optimal policy? (Sec. 4.4)
  • Is S-ICQL robust to the quality of offline datasets and hyperparameters? (Sec. 4.5 and Appendix G)
  • Can S-ICQL encode task-relevant information for efficient prompt construction? (Sec. 4.6)
Environments. We conduct experiments on three challenging benchmarks commonly used for evaluating ICRL algorithms: (i) DarkRoom Laskin et al. (2023), a 2D discrete environment where the agent must locate an unknown goal. (ii) MuJoCo Todorov et al. (2012), a standard testbed including tasks with varying reward functions and transition dynamics. (iii) Meta-World ML1 Yu et al. (2020), a robotic benchmark with 50 manipulation tasks. Tasks are randomly sampled and split into training sets $M^{\text{train}}$ and test sets $M^{\text{test}}$. Further environment details are provided in Appendix C.
Pretraining Datasets. For DarkRoom, suboptimal datasets are generated using a noisy action selection strategy by combining optimal and random policies. For MuJoCo and Meta-World, datasets are collected using a single task RL policy for each task. We construct two qualities of offline datasets: Mixed and Medium. More details on dataset construction are provided in Appendix D.
Baselines. We compare to six competitive ICRL approaches and one offline meta-RL method, including: 1) IC-IQL (Tarasov et al., 2025), 2) DIT (Dong et al., 2024), 3) DICP (Son et al., 2025), 4) IDT (Huang et al., 2024), 5) AD (Laskin et al., 2023), 6) DPT (Lee et al., 2023), and 7) UNICORN (Li et al., 2024b). Details about these baselines are provided in Appendix E.
To ensure a fair comparison, we conduct a few-shot evaluation for all methods. During testing, each policy directly interacts with the environment for a few episodes using fixed parameters, conditioned on a prompt sampled from past interactions. Results are reported as the mean of 10 trials with 95% bootstrapped confidence intervals, and standard errors are also provided.
Figure 2: Few-shot evaluation return curves of S-ICQL and baselines on Mixed datasets.
Table 1: Few-shot evaluation returns of S-ICQL and baselines on Mixed datasets, i.e., numerical results of converged performance from Figure 2.
@@ 4.1 Main Results
The baselines typically work under the few-shot setting, since they rely on task prompts or warm-start data to infer task representations. We evaluate S-ICQL against these baselines in an aligned few-shot setting, where all methods utilize the same number of interaction trajectories for task inference. Figure 2 illustrates the evaluation return curves across various environments using Mixed datasets, and Table 1 summarizes the numerical results of converged performance. Across environments with varying reward functions and transition dynamics, S-ICQL consistently demonstrates superior data efficiency and higher asymptotic performance. In more complex environments such as Ant-Dir and HalfCheetah-Vel, the advantage of S-ICQL is more pronounced, highlighting its promising in-context learning abilities on challenging tasks. Moreover, S-ICQL generally exhibits lower variance during the testing phase, indicating both improved data efficiency and better training stability.
@@
Table 2: Few-shot converged returns of S-ICQL and its ablations on Mixed datasets.
Ablation Reach Cheetah-Vel Ant-Dir
w/o_cq 736.72 $\pm$ 39.42 -78.35 $\pm$ 4.50 591.31 $\pm$ 45.54
w/o_c 792.09 $\pm$ 6.66 -56.19 $\pm$ 1.94 693.87 $\pm$ 30.41
w/o_q 752.41 $\pm$ 24.15 -63.66 $\pm$ 8.26 784.07 $\pm$ 24.88
S-ICQL 806.97 $\pm$ 6.35 -35.48 $\pm$ 1.33 813.34 $\pm$ 14.12
@@ 4.2 Ablation Study
To assess the respective contribution of each component, we compare S-ICQL with three ablations: (i) w/o_c, which removes the world modeling component and directly uses a trajectory of raw transitions to construct prompts; (ii) w/o_q, which removes the Q-learning component and learn in-context policies in a pure supervised paradigm; and (iii) w/o_cq, which removes both world modeling and Q-learning, reducing the model to the original DPT. In all ablations, the remaining structural components are kept identical to those in the full S-ICQL.
Table 2 present ablation results across representative environments (refer to Appendix H for graphical results). First, ablating the world model causes a notable performance drop, especially in complex tasks like Ant-Dir, emphasizing its importance for precise in-context inference. Second, removing Q-learning reduces performance, underscoring its role in refining policies using suboptimal data. Integrating Q-learning enables S-ICQL to improve policies from noisy or inferior trajectories, enhancing its stitching ability to maximize rewards. Finally, removing both components results in the greatest decline, reducing the model to simple behavioral cloning that fails to generalize in complex tasks. Overall, the ablation results validate S-ICQL’s effectiveness in harnessing both world modeling and Q-learning for efficient reward maximization and generalization across tasks.
Table 3: Few-shot evaluation returns of S-ICQL and baselines for OOD tasks on Mixed datasets, i.e., numerical results of converged performance from Figure 3.
Method UNICORN DPT AD IDT DICP DIT IC-IQL S-ICQL
[…]
4.3 Generalization to Out-of-Distribution (OOD) Tasks
Figure 3: Few-shot evaluation curves of S-ICQL and baselines for OOD tasks on Mixed datasets.
Acknowledging the importance of evaluating ICRL algorithms under distribution shifts, we test S-ICQL against all baselines on OOD tasks across representative environments. The target velocities for Cheetah-Vel range from $[0.1,3.0]$, and the target directions for Ant-Dir range from $0,2\pi]$. We split tasks by target values to construct an OOD setting. The model is trained on tasks with a lower range of target velocities/directions and tested on the remaining higher range, ensuring that training and testing tasks come from different distributions. This setup allows us to assess S-ICQL’s ability to generalize beyond the set of training tasks. As shown in Figure 3 and Table 3, S-ICQL consistently outperforms baselines on OOD tasks. We extrapolate the meta-level knowledge across tasks by the extrapolation ability of the world model, which is more accurate and robust since the world model is intrinsically invariant to behavior policies or collected datasets. The world model shares some common structure across the task distribution (even for OOD tasks), e.g., the kinematics principle or locomotion skills. Hence, the extrapolation of the world model also works for OOD test tasks in this case. This observation is also consistent with the visualization results in Sec. 4.6.
4.4 Stitching Capability
Figure 4: Comparison of best dataset returns with DPT and S-ICQL performances on training tasks.
We conducted an in-depth analysis to examine whether S-ICQL truly exhibits stitching capability. Specifically, we pretrained the model on offline datasets containing only suboptimal trajectories and evaluated it on the corresponding training tasks. Figure 4 presents a comparison between the best returns in the offline datasets (denoted as OPT for optimal dataset return)and the performance achieved by DPT, w/o_q (S-ICQL without Q-learning), and S-ICQL, including both the evaluation return curves and the final return values. As w/o_q and DPT follows the supervised pretraining paradigm, its performance never exceeds the dataset best, underscoring the limits of pure supervised methods. In contrast, S-ICQL achieves performance exceeding the best returns in the dataset on training tasks, providing strong evidence of genuine stitching capability by composing suboptimal trajectory segments into globally optimal policies.
@@ 4.5 Robustness to the Quality of Offline Datasets
Table 4: Few-shot converged results of S-ICQL and baselines on Medium datasets.
@@
[…]
To evaluate S-ICQL’s robustness to data quality, we conduct experiments on Medium datasets containing only suboptimal data. As shown in Figure 5 and Table 4, S-ICQL outperforms all baselines, particularly in complex environments like Cheetah-Vel and Ant-Dir, showing its superiority when learning from suboptimal data. Notably, many baselines converge to suboptimal policies using imperfect data, especially for pure imitation learning methods like DPT. This again highlights our advantage of using Q-learning to not only simply imitate collected data, but also to combine parts of suboptimal trajectories for finding globally optimal behaviors. These results, along with those in Sec. 4.1, confirm the robustness of our method to varying dataset quality.
Figure 5: Few-shot evaluation return curves of S-ICQL and baselines on Medium datasets.